In need of help!!!!
This spinner is spun 36 times. The spinner landed on A 6 times, on B 21 times, and on C 9 times. Compute the empirical probability that the spinner will land on B.

Found 2 solutions by jim_thompson5910, stanbon:
Answer by jim_thompson5910(35256)

P(B) = (number of times spinner landed on B)/(number of times spinner was spun)

P(B) = 21/36

P(B) = 7/12

Answer by stanbon(75887)

This spinner is spun 36 times. The spinner landed on A 6 times, on B 21 times, and on C 9 times. Compute the empirical probability that the spinner will land on B.
----
Empirical means "based on your experience".
---
Answer: 21/36 = 7/12
==========================
Cheers,
Stan H.